An ostrich egg has about the same volume as a sphere with a diameter of 5 inches. If the eggshell is 1/12 inch thick, find the volume of just the shell, NOT including the interior of the egg. Round to the nearest tenth.
Can you please show the steps too, please? Thanks!!!

Answer by rothauserc(4718)

the formula for the volume(V) of a sphere is
V = (4/3)*pi*r^3
********************************************
volume of the eggshell, note that (4 11/12) / 2 = 2.46
V = (4/3)*pi*(2.5)^3 - (4/3)*pi*(2.46)^3
V = 65.449846958 - 62.358251704
V = 3.091595254
volume of the eggshell is 3.09 cubic inches